[Hallc_running] [New Logentry] Hall C RC update: Thursday, Dec. 16, 2021

gaskelld at jlab.org gaskelld at jlab.org
Thu Dec 16 08:30:02 EST 2021

Logentry Text:
Yesterday at 4pm, we completed running on the SHMS right (Q2=6, W=2.4, low epsilon) setting, and rotated the SHMS to 11.12 degrees for the SHMS center setting.   Total good pion events on LH2 at the previous setting were 16705 (not quite matching the LD2 statistics, 17797).

There were a few RF related issues overnight, leading to slightly lower efficiency on owl shift, but we have accumulated 9296 good pion events on LH2 as of 7:30 this morning.  We will discuss the timing of the switch to the aluminum dummy target at the 10 am RC meeting this morning.

RC meeting bluejeans link:

RC meeting page:


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3964146
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211216/aa8671dc/attachment.html>

More information about the Hallc_running mailing list